# Movement Prediction-Adjusted Naive Forecast: Is the Naive Baseline Unbeatable in Financial Time Series Forecasting?

### v10

@@ 3 Experiments
This section details the experiments for evaluating the performance of the MPANF using real world data. We first present the experimental setup, including the characteristics of the datasets, the baselines, and the evaluation metrics. We subsequently present and analyze the experimental results, which cover a comparative assessment of out-of-sample performance, and a validation of the retrospective condition.
@@ 3.1.1 Datasets and Preprocessing
Because the MPANF’s effectiveness depends on the quality of the given movement prediction, data selection should account for both the target series and the corresponding movement prediction availability. Given the well-documented strong co-movement between the Financial Times Stock Exchange (FTSE) index and U.S. stock prices (Sarwar,, 2020; Gupta et al.,, 2024), together with the fact that the FTSE open price is recorded approximately 13 hours before the U.S. market close price on the same date, we selected eight U.S. financial time series as the target series, including Apple (AAPL), Boeing (BA), Google (GOOGL), Goldman Sachs (GS), Halliburton (HAL), Brent crude oil (OIL), Schlumberger (SLB), and Tesla (TSLA), on the basis of their strong predictive alignment with FTSE movements. Accordingly, the FTSE open price is used as the exogenous variable for generating movement predictions. The details of the movement prediction method using an exogenous variable are provided in Appendix C.
Historical data for the FTSE open price and the close prices of the eight target series were obtained from Yahoo Finance, covering the period from October 8, 2014, to September 13, 2024. The FTSE data were aligned with each target series on a day-by-day basis, with extra observations removed and missing values imputed via forward filling (Junninen et al.,, 2004). After data alignment, all time series were standardized to a uniform length of 2,500 observations and then split into in-sample and out-of-sample sets at a 1:1 ratio, ensuring sufficient data for both model estimation and evaluation. The close prices of the eight target variables and the FTSE open price are illustrated in Figures 2 and 3, respectively, with a brief description provided in Table 1. All eight target series achieved in-sample movement prediction accuracies above 0.55, suggesting that the directional signals from the FTSE opening price carried useful predictive information. For each target series, the out-of-sample forecasts by the MPANF were produced via Algorithm 1.
@@ 3.1.2 Baselines and Evaluation Metrics
As improvements over naive forecasts in financial time series forecasting can be difficult to achieve, even with sophisticated approaches such as machine learning and deep learning, we adopt simple and widely recognized baselines, including the naive forecast, naive forecast with drift, IMA(1,1), and linear regression, for performance comparison Rogers, (1995); Jadevicius and Huston, (2015); Franses, (2020). In particular, the linear regression model, which acts as a meta-learner to combine the naive forecast and the movement prediction, uses only the current observed value and the predicted movement as predictors.
@@ 3.2.1 Out-of-Sample Performance
The out-of-sample forecasting results for the eight financial time series are illustrated in Figure 4. Given that each out-of-sample set consists of 1250 time steps and that predictions from different methods visually overlap to a large extent, only the first five predicted points from each set are displayed for clarity. Small differences between the MPANF and the other forecasting methods are observed because the overall in-sample movement prediction accuracy for each time series is just slightly above 0.55, leading to a small coefficient $\alpha$, which means that the magnitude of adjustment for each time step is approximately one tenth of the base magnitude. The naive forecast, naive forecast with drift, and IMA(1,1) forecasts yield similar predictions, reinforcing the suitability of these methods as baselines and the robustness of the naive forecast as the benchmark for time series forecasting. Moreover, because the linear regression model closely mirrors the structure of the MPANF, its performance is similarly dependent on movement prediction accuracy and therefore remains close to that of the MPANF.
@@
Tables 2, 3, 4, and 5 present performance comparisons between the MPANF and the baseline methods for eight out-of-sample sets, reported in terms of the RMSE, MAE, MAPE, and sMAPE, respectively. Notably, the MPANF consistently outperforms the naive forecast, with the only exceptions being the MAPE and sMAPE on the GOOGL dataset. Specifically, the MPANF achieves the lowest RMSE in 4 out of 8 cases, the lowest MAE in 6 out of 8 cases, the lowest MAPE in 6 out of 8 cases, and the lowest sMAPE in 5 out of 8 cases, demonstrating its overall superiority over the baseline methods.
@@
In some cases, baseline methods such as naive forecasting with drift, IMA(1,1), or linear regression have demonstrated superior performance, yielding lower error metric values than other models do. For example, these methods produced lower RMSE values for AAPL, GOOGL, HAL, and OIL, and lower MAEs, MAPEs, and sMAPEs for AAPL and GOOGL. However, the performances of these methods are inconsistent; they frequently fall short of the naive forecast in other cases.
Compared with the naive forecast, linear regression may sometimes perform better, particularly when movement prediction is used as an input feature. However, this occasional improvement lacks a clear theoretical justification. Its performance generally falls short of the MPANF because of a higher susceptibility to overfitting. While linear regression can fit in-sample data closely, its out-of-sample performance degrades significantly if the accuracy of the movement prediction decreases. In contrast, the MPANF’s coefficient determination is based on theoretically derived optimal conditions, rather than numerical minimization of in-sample errors. This theoretical guidance acts as a built-in regularizer, which effectively mitigates overfitting and improves the method’s ability to generalize to new and unseen data.
As shown in Figures 5, 6, 7, and 8, the MPANF outperforms the baseline methods in both accuracy and robustness across the four evaluation metrics. These empirical results align with our theoretical analysis and confirm the MPANF’s consistent improvement over the naive forecast. This superior performance, even in challenging financial scenarios characterized by heteroscedasticity and high volatility, highlights the method’s broad and reliable applicability.
@@ 3.2.2 Retrospective Validation
Table 6 summarizes the validation results for each target series. Across the eight series that exhibited improved out-of-sample RMSEs, the retrospective condition was met in all the cases. Moreover, the coefficient between the in-sample and out-of-sample periods decreased in six out of eight cases, confirming the robustness of the MPANF even under the decline in movement prediction accuracies for out-of-sample forecasting.
@@ 4 Discussion
The implications of the findings and the potential limitations of the study are summarized in four points as follows.
First, in many real-world forecasting applications, decision-making requires the integration of different types of forecasts, such as binary labels and continuous values. Such integration can be technically challenging. The MPANF offers a practical solution by converting binary movement predictions into continuous point forecasts, thereby facilitating more informed decisions in diverse forecasting scenarios. The formulation of the MPANF also implies the value of incorporating directional information into a naive forecast and suggests that, in addition to developing complex regression models, we can instead develop a movement prediction-based model for accurate point forecasting.
Second, the coefficient or scaling factor in the MPANF is derived analytically on the basis of the assumption that the movement prediction is unbiased. However, although the MPANF imposes no restrictions on the choice of movement prediction models, this assumption can be violated in practice, especially when the movement prediction models that are prone to overfitting suffer substantial degradation in out-of-sample performance. Such a violation could lead to a substantial gap between the analytical estimate and the true optimum for the coefficient, reducing the practical advantages of the MPANF over the naive baseline. Moreover, generating consistent and reliable movement predictions across different segments of a financial time series remains a challenging task. Consequently, we cannot define a threshold of movement prediction accuracy for the effectiveness of MPANF, and its applicability is conditional on the availability of a meaningful and robust movement prediction for the target variable.
Third, the retrospective condition formalized in Proposition 3 offers more than a binary assessment of outperformance; it provides meaningful interpretability and practical insights into the method’s behavior. Specifically, it reveals that the MPANF’s predictive gains over the naive baseline occur only when the product of its out-of-sample movement prediction accuracy and the mean absolute increment is sufficiently large, highlighting the crucial interplay between movement prediction and increment magnitude for point forecasting performance. In addition, it serves as a valuable diagnostic tool: if the MPANF fails to improve upon the naive baseline, one can examine whether the Eq. (15) was violated, thereby distinguishing between method inadequacy and inherent unpredictability in the out-of-sample period. Furthermore, it encourages practices that enhance the stability of movement prediction by suggesting that optimal performance is approached when in-sample estimates align closely with their out-of-sample counterparts. Thus, while the condition is retrospective in nature, its implications are forward-looking, informing both interpretation and future method tuning.
Finally, while the empirical results show that even relatively modest directional signals can enhance point forecasts, the magnitude of such improvements is often marginal and context dependent. These improvements do not necessarily translate into economic benefits, especially when real-world factors such as transaction costs, slippage, and market frictions are taken into account. Trading decisions can be made after the MPANF’s performance under higher movement prediction accuracies is examined.
@@ 5 Conclusions
This study introduces the MPANF, a forecast combination method that adjusts the naive forecast by incorporating the movement prediction of the target variable. The proposed method demonstrates that systematic improvement over the naive baseline becomes theoretically and empirically possible when movement predictions are accurate and reliable. Under these conditions, movement prediction introduces genuine directional information that the naive forecast alone cannot provide. Thus, MPANF reframes the long-standing belief in the naive baseline’s supremacy: its apparent "unbeatability" in financial time series forecasting is not due to inherent optimality, but rather a consequence of the scarcity of reliable directional information. The naive forecast’s dominance, therefore, is a conditional property, not a universal one.
Several avenues for future research emerge from this work. First, because the present evaluation was conducted under modest directional accuracy, developing more advanced movement prediction models and assessing MPANF performance at higher accuracy levels could enhance its practical value in real-world trading. Second, applying MPANF to other time series domains, particularly those where the naive forecast is hard to beat yet movement prediction is feasible, could broaden its applicability. Third, extending the MPANF beyond the univariate case would address the many forecasting tasks that arise in high-dimensional settings. Finally, while this study employs a fixed adjustment for simplicity, exploring adaptive adjustment schemes may further improve the forecasting accuracy.
@@ Appendix A. Algebraic Derivation of $\Delta MSE_{\text{in}}$
$MSE_{\text{in}}^{\text{Naive}}=\frac{1}{N}\left(\sum_{t\in\text{in}}\left(y_{t}-y_{t-1}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{in}}\left((y_{t-1}+d_{t}\cdot|\epsilon_{t}|)-y_{t-1}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$ (A1)
@@
$MSE_{\text{in}}^{\text{MPANF}}=\frac{1}{N}\left(\sum_{t\in\text{in}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(y_{t}-\hat{y}_{t}\right)^{2}+\sum_{t\in\text{incorrect}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left[(y_{t-1}+d_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{d}_{t}\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}})\right]^{2}\right$.
@@
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$ (A2)
@@
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)-\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right$.
@@
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}-\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)-\sum_{t\in\text{incorrect}}\left(2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$ (A3)
@@
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in}}$ (A4)
@@
$\approx\frac{1}{N}\left(2\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}}\cdot(n_{\text{in}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in}})-n_{\text{in}}^{\text{correct}}\cdot\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}-2\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}}\cdot(n_{\text{in}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in}})-n_{\text{in}}^{\text{incorrect}}\cdot\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)$

### v11

@@ 3 Experiments
This section details the experiments for evaluating the performance of the MPANF using real world data. We first present the experimental setup, including the characteristics of the datasets, the baselines, and the evaluation metrics. Subsequently, we present and analyze the experimental results, which include a comparative assessment of out-of-sample performance, and a validation of the retrospective condition.
@@ 3.1.1 Datasets and Preprocessing
Because the MPANF’s effectiveness depends on the quality of the given movement prediction, data selection should account for both the target series and the corresponding movement prediction availability. Given the well-documented strong co-movement between the Financial Times Stock Exchange (FTSE) index and U.S. stock prices (Sarwar, 2020; Gupta et al., 2024), together with the fact that the FTSE open price is recorded approximately 13 hours before the U.S. market close price on the same date, we selected eight U.S. financial time series as the target series, including Apple (AAPL), Boeing (BA), Google (GOOGL), Goldman Sachs (GS), Halliburton (HAL), Brent crude oil (OIL), Schlumberger (SLB), and Tesla (TSLA), on the basis of their strong predictive alignment with FTSE movements. Accordingly, the FTSE open price is used as the exogenous variable for generating movement predictions. The details of the movement prediction method, which is based on the co-movement of the target variable and an exogenous variable, are provided in Appendix C.
Historical data for the FTSE open price and the close prices of the eight target series were obtained from Yahoo Finance, covering the period from October 8, 2014, to September 13, 2024. The FTSE data were aligned with each target series on a day-by-day basis, with extra observations removed and missing values imputed via forward filling (Junninen et al., 2004). After data alignment, all time series were truncated to a uniform length of 2,500 observations and then split into in-sample and out-of-sample sets at a 1:1 ratio, ensuring sufficient data for both model estimation and evaluation. The close prices of the eight target variables and the FTSE open price are illustrated in Figures 2 and 3, respectively, with a brief description provided in Table 1. Using the movement of the FTSE open price as the movement prediction for all eight target series, the in-sample movement prediction accuracy for each dataset exceeded 0.55, indicating that the directional signals derived from the FTSE open price are meaningful. For each target series, the out-of-sample forecasts by the MPANF were produced via Algorithm 1.
@@ 3.1.2 Baselines and Evaluation Metrics
As improvements over naive forecasts in financial time series forecasting can be difficult to achieve, even with sophisticated approaches such as machine learning and deep learning, we adopt simple and widely recognized baselines, including the naive forecast, naive forecast with drift, IMA(1,1), and linear regression, for performance comparison Rogers (1995); Jadevicius and Huston (2015); Franses (2020). In particular, the linear regression model, which acts as a meta-learner to combine the naive forecast and the movement prediction, uses only the current observed value and the predicted movement as predictors.
@@ 3.2.1 Out-of-Sample Performance
The out-of-sample forecasting results for the eight financial time series are illustrated in Figure 4. Given that each out-of-sample set consists of 1250 time steps and that predictions from different baseline methods visually overlap to a large extent, only the first five predicted points from each set are displayed for clarity. Small differences between the MPANF and the baseline methods are observed because the overall in-sample movement prediction accuracy for each time series is just slightly above 0.55, leading to a small coefficient $\alpha$, which means that the magnitude of adjustment for each time step is approximately one tenth of the base magnitude. Moreover, the naive forecast, naive forecast with drift, and IMA(1,1) forecasts yield similar predictions, reinforcing the suitability of these methods as baselines and the robustness of the naive forecast as the benchmark for time series forecasting. Moreover, because the linear regression model closely mirrors the structure of the MPANF, its performance is similarly dependent on movement prediction accuracy and therefore remains close to that of the MPANF.
@@
Tables 2, 3, 4, and 5 present performance comparisons between the MPANF and the baseline methods for eight out-of-sample sets, reported in terms of the RMSE, MAE, MAPE, and sMAPE, respectively. Notably, the MPANF consistently outperforms the naive forecast across all four metrics, except for the MAPE and sMAPE on the GOOGL dataset, where its performance is slightly inferior. Specifically, the MPANF achieves the lowest RMSE in 4 out of 8 cases, the lowest MAE in 6 out of 8 cases, the lowest MAPE in 6 out of 8 cases, and the lowest sMAPE in 5 out of 8 cases, demonstrating its overall superiority over the baseline methods. On average, the MPANF achieved the following improvements over the simple baseline model: a 0.32% reduction in the RMSE, a 0.41% reduction in the MAE, a 0.47% reduction in the MAPE, and a 0.46% reduction in the sMAPE.
@@
In addition, baseline methods such as the naive forecast with drift, IMA(1,1), or linear regression have demonstrated superior performance in certain time series, yielding lower error metric values than other methods do. For example, they produced lower RMSE values for AAPL, GOOGL, HAL, and OIL, and lower MAEs, MAPEs, and sMAPEs for AAPL and GOOGL. However, the performances of these methods are inconsistent; they frequently fall short of the naive forecasts in other time series.
Compared with the naive forecast, linear regression can sometimes perform better, particularly when movement prediction is used as an input feature. However, this occasional improvement lacks a clear theoretical justification, and its performance generally falls short of the MPANF because of a higher susceptibility to overfitting. While linear regression can fit in-sample data closely, its out-of-sample performance degrades significantly if the accuracy of the movement prediction decreases. In contrast, the coefficient in the MPANF is analytically derived rather than obtained through numerical minimization of in-sample errors. As an analytical estimate of the optimal value, the coefficient $\hat{\alpha}_{\text{in}}^{*}$ may deviate slightly from the true empirical optimum. This controlled deviation constrains the model’s flexibility, thereby reducing the risk of overfitting to in-sample noise and enhancing the stability of out-of-sample performance.
As shown in Figures 5, 6, 7, and 8, the MPANF outperforms the baseline methods across all four evaluation metrics. These empirical results align with our theoretical analysis and confirm the MPANF’s consistent improvement over the naive baseline. This superior performance, even in challenging financial scenarios characterized by heteroscedasticity and high volatility, highlights the method’s broad and reliable applicability.
@@ 3.2.2 Retrospective Validation
Table 6 summarizes the validation results for each target series. Among the eight series where the MPANF achieved lower out-of-sample RMSE than the naive baseline, the retrospective condition held in all cases. Moreover, the in-sample coefficient $\hat{\alpha}^{*}_{\text{in}}$ is smaller than its out-of-sample counterpart $\hat{\alpha}^{*}_{\text{out}}$ in six of the eight series, indicating the robustness of the proposed method even when the movement-prediction accuracy decreased at the out of sample period.
@@ 4 Discussion
This section discusses the implications of the findings and the potential limitations of the study. We begin with the mechanism, interpretability, and potential value of the proposed method and then detail the conditions that constrain its performance.
Many real-world decision-making processes require combining heterogeneous forecasts such as binary labels and continuous values. Under this condition, MPANF provides a practical combination approach by translating movement predictions into point forecasts, enabling more informed decisions when directional signals are available and reliable. Its coefficient, which is a scaling factor, is derived in closed form under an unbiased movement prediction assumption. This analytical estimate can be regarded as a stable, regularized weight, which avoids noisy numerical tuning and makes the combination step lightweight, interpretable, and easy to deploy. Proposition 3 further states that MPANF improves on the naive baseline only if the out-of-sample product $\hat{\alpha}_{\text{out}}^{*}\cdot\bar{\epsilon}_{\text{out}}$ is at least half of the in-sample product $\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}$. In short, MPANF offers a clear mechanism for turning directional information into statistical gains, and it indicates a pathway via stronger movement prediction for potential economic relevance.
However, the effectiveness of MPANF is bottlenecked by the strength, bias, and stability of the directional signal. When directional signals are weak or unstable, MPANF may still underperform the naive baseline. In particular, if the assumption of an unbiased movement prediction does not hold, a substantial gap can arise between the analytical estimate and the empirical optimum, further diminishing the gains over the naive baseline. Absence of improvement should be diagnosed with Eq. (15) to distinguish method limitations from genuine unpredictability and to guide remedies, such as developing more advanced movement prediction models. Finally, modest statistical gains may not translate into economic value after taking into account the transaction costs, slippage, and market frictions. Therefore, practical impact ultimately depends on directional signals that are both accurate and robust.
@@ 5 Conclusions
This study introduces the MPANF, a forecast combination method that adjusts the naive forecast by incorporating the movement prediction of the target variable. The proposed method demonstrates that systematic improvement over the naive baseline becomes theoretically and empirically possible when movement predictions are accurate and reliable. Thus, MPANF reframes the long-standing belief in the naive baseline’s supremacy: its apparent ”unbeatability” in financial time series forecasting is not due to inherent optimality, but rather a consequence of the scarcity of reliable directional information. The dominance of the naive baseline, therefore, is a conditional property, not a universal one.
This study offers several directions for future research. First, since the current experiments employed movement predictions with modest accuracies, developing more advanced movement prediction models and evaluating the MPANF at higher accuracy levels would further demonstrate its effectiveness and enhance its practical value in real trading scenarios. Second, applying the MPANF to other types of time series beyond financial data, particularly those where the naive baseline is difficult to surpass but movement prediction is feasible, would broaden the applicability of the proposed method. Third, extending the MPANF to handle multivariate time series would enable testing its effectiveness in more complex, high-dimensional forecasting settings. Finally, while this study used a fixed adjustment factor for simplicity, developing adaptive adjustment schemes could further improve the method’s performance.
@@ Appendix A. Algebraic Derivation of $\Delta MSE_{\text{in}}$
$MSE_{\text{in}}^{\text{Naive}}=\frac{1}{M}\left(\sum_{t\in\text{in}}\left(y_{t}-y_{t-1}\right)^{2}\right)$
$=\frac{1}{M}\left(\sum_{t\in\text{in}}\left((y_{t-1}+d_{t}\cdot|\epsilon_{t}|)-y_{t-1}\right)^{2}\right)$
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$ (A1)
@@
$MSE_{\text{in}}^{\text{MPANF}}=\frac{1}{M}\left(\sum_{t\in\text{in}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(y_{t}-\hat{y}_{t}\right)^{2}+\sum_{t\in\text{incorrect}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left[(y_{t-1}+d_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{d}_{t}\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}})\right]^{2}\right$.
@@
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right$.
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$ (A2)
@@
$\quad-\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right$.
@@
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}-\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)-\sum_{t\in\text{incorrect}}\left(2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$ (A3)
@@
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{in}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{in}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in}}$ (A4)
@@
$\approx\frac{1}{M}\left(2\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}}\cdot(m_{\text{in}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in}})-m_{\text{in}}^{\text{correct}}\cdot\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right$.
$\quad\left.-2\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}}\cdot(m_{\text{in}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in}})-m_{\text{in}}^{\text{incorrect}}\cdot\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)$